# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 2.4. Haar measures
We fix $\eta_{{\bf K}_{\infty}}$ so that the measure of ${\bf K}_{\infty}$ is $1$. Then we normalize $\eta_{{\bf Sp}_{2}({\mathbb{R}}))}$ in such a way that the quotient $\eta_{{\bf Sp}_{2}({\mathbb{R}})}/\eta_{{\bf K}_{\infty}}$ corresponds to the measure $(\det Z)^{-3}{\rm{d}}X{\rm{d}}Y$ on the Siegel upper-half space ${\mathfrak{h}}_{2}\cong{\bf Sp}_{2}({\mathbb{R}})/{\bf K}_{\infty}$. By $\mathsf{G}({\mathbb{R}})^{0}={\mathsf{Z}}({\mathbb{R}})^{0}\,{\bf Sp}_{2}({\mathbb{R}})\cong{\mathbb{R}}_{>0}\times{\bf Sp}_{2}({\mathbb{R}})$, $\eta_{\mathsf{G}({\mathbb{R}})}$ is defined by demanding that its restriction to $\mathsf{G}({\mathbb{R}})^{0}$ is $\eta_{{\mathbb{R}}^{\times}}\otimes\eta_{{\bf Sp}_{2}({\mathbb{R}})}$. For $p<\infty$, we fix $\eta_{\mathsf{G}({\mathbb{Q}}_{p})}$ by ${\operatorname{vol}}({\bf K}_{p})=1$. Then, $\eta_{\mathsf{G}({\mathbb{A}})}$ is defined to be the restricted product of $\eta_{\mathsf{G}({\mathbb{Q}}_{v})}\,(v\leqslant\infty)$.
2.5. Idele class characters
Let $\Lambda$ denote a character of the finite group
${\rm Cl}(E):={\mathbb{A}}_{E}^{\times}/E^{\times}E_{\mathbb{R}}^{\times}\widehat{\mathfrak{o}}_{E}^{\times}.$
As is well known, this group is isomorphic to the ideal class group of $E$. We regard $\Lambda$ as an idele class character of $E^{\times}$ of finite order. Since $^{\mathbb{A}}^{\times}={\mathbb{Q}}^{\times}({\mathbb{R}}_{>0})\widehat{\mathbb{Z}}^{\times}$ is contained in the subgroup $E^{\times}E_{\mathbb{R}}^{\times}\widehat{\mathfrak{o}}_{E}^{\times}$, we have
(2.23) $\Lambda|{\mathbb{A}}^{\times}={\bf 1}.$
Let $v$ be a place of $E$ and $\Lambda_{v}$ the $v$-component of $\Lambda$, i.e., $\Lambda=\otimes_{v}\Lambda_{v}$. We say that $v$ is inert in $E/{\mathbb{Q}}$, splits in $E/{\mathbb{Q}}$ or ramifies in $E/{\mathbb{Q}}$ if $E_{v}:=E\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ is an unramified field extension of ${\mathbb{Q}}_{p}$, is isomorphic to ${\mathbb{Q}}_{p}\oplus{\mathbb{Q}}_{p}$, or is a ramified field extension of ${\mathbb{Q}}_{p}$, respectively. Since $\Lambda$ is supposed to be trivial in $E_{\mathbb{R}}^{\times}$, we have $\Lambda_{\infty}={\bf 1}$; moreover,
(2.24) If v is inert in E/, then $\text{If $v$ is inert in $E/{\mathbb{Q}}$, then $\Lambda_{v}={\bf 1}$}.$
Indeed, since $\Lambda$ is supposed to be trivial on $\widehat{\mathfrak{o}}_{E}^{\times}$, the restriction $\Lambda_{v}|{\mathfrak{o}_{E,v}^{\times}}$ is trivial. If $v$ is inert and $p$ denote the residue characteristic of $E_{v}$, then $p\in{\mathbb{Q}}_{p}$ is a prime element of the local field $E_{v}$. We have $\Lambda_{v}(p)=1$ due to (2.23). Hence, $\Lambda_{v}$ is trivial on $p^{{\mathbb{Z}}}\mathfrak{o}_{E,v}^{\times}=E_{p}^{\times}$.
Define the Galois conjugate $\Lambda^{\dagger}$ of $\Lambda$ by setting
@@ 2.5. Idele class characters
We have $\Lambda(\tau)\Lambda^{\dagger}(\tau)=\Lambda(\tau\bar{\tau})=1$ for $\tau\in{\mathbb{A}}_{E}^{\times}$ by (2.23). Hence
@@
For an idele class character $\xi$ of $E^{\times}$, let $\widehat{L}(s,\xi)$ be the completed Hecke $L$-function of $\xi$, and $L_{p}(s,\xi)$ its local $p$-factor for $p\leqslant\infty$. Then, $L_{p}(s,\xi^{\dagger})=L_{p}(s,\xi)$ for any $p<\infty$. Indeed, if $p$ is not ramified in $E/{\mathbb{Q}}$, the equality is trivial. Suppose that $E/{\mathbb{Q}}$ is ramified; if $\xi_{p}$ is a ramified character of $E_{p}^{\times}$, then both $L$-factors are $1$. If $\xi_{p}$ is unramified, then for any prime element $\varpi$ of $E_{p}$, we have $(\xi^{\dagger})_{p}(\varpi)=\xi_{p}(\bar{\varpi})=\xi_{p}(\varpi)$, which implies the equality between local $p$-factors.
For any character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}^{\times}\rightarrow{\mathbb{C}}^{1}$, we define $\mu_{E}:=\mu\circ{\operatorname{N}}_{E/{\mathbb{Q}}}:{\mathbb{A}}^{\times}_{E}/E^{\times}\rightarrow{\mathbb{C}}^{1}$. Then, $\mu_{E}$ is Galois invariant, and $\mu_{E}|E_{\infty}^{\times}={\bf 1}$. Hence, $(\Lambda\mu_{E})^{\dagger}=\Lambda^{\dagger}\mu_{E}^{\dagger}=\Lambda^{-1}\mu_{E}$, and
@@ 3.1. Eisenstein series
Let $\mathcal{S}({\mathbb{A}}_{E}^{2})$ be the space of all Schwartz-Buruhat functions on ${\mathbb{A}}_{E}^{2}$. The Fourier transform $\widehat{\phi}$ of $\phi\in\mathcal{S}({\mathbb{A}}_{E}^{2})$ is defined by
@@
with $c_{p}=1$ if $p<\infty$ and $c_{\infty}=\frac{\sqrt{|D|}}{2}$ if $p=\infty$, where $\varepsilon(s,\bullet,\psi_{E_{p}})$ denote Tate’s local epsilon factor defined by the character $\psi_{E_{p}}$ of $E_{p}$ and the associated self-dual measure on $E_{p}$ as usual. For a decomposable element $\phi=\otimes_{p{\leqslant}\infty}\phi_{p}\in\mathcal{S}({\mathbb{A}}_{E}^{2})$, we define
@@
Note that $\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}$ is left ${\mathsf{B}}^{\#}({\mathbb{Q}})$- invariant and right ${\bf K}^{\#}$-finite.
The Eisenstein series attached to $\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}$ is defined by the absolutely convergent series
@@
for ${\operatorname{Re}}(s)>1$. The following property of the Eisenstein series is standard.
@@ 3.2. Rankin-Selberg integral and the basic identity
By Lemma 3.1(iii) and by the Fourier expansion (4.5), it is straightforward to show the absolute convergence of the integral for $s\in{\mathbb{C}}$ where the Eisenstein series is regular. Let $\psi:{\mathbb{A}}/{\mathbb{Q}}\rightarrow{\mathbb{C}}^{1}$ be the basic character. For $T\in{\mathsf{V}}({\mathbb{Q}})$, define a character of $\mathsf{N}({\mathbb{A}})$ by
@@
The $(T_{\theta},\Lambda)$-Bessel period of a cusp form $\varphi$ on ${\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})$ is defined by the integral111 Note that $\mathsf{m}(I_{\theta}(a),{\operatorname{N}}_{E/{\mathbb{Q}}}(a))=a1_{4}$ for $a\in{\mathbb{A}}^{\times}$. Then, due to $\varphi$ being ${\mathsf{Z}}({\mathbb{A}})$-invariant, the integral $B^{T_{\theta},\Lambda}(\varphi;g)$ is 00 if $\Lambda|{\mathbb{A}}^{\times}\not=1$. This trivial vanishing is not the case by (2.23).
@@
for $g\in\mathsf{G}({\mathbb{A}})$, where ${\rm{d}}^{\times}\tau$ denote the quotient measure on ${\mathbb{A}}_{E}^{\times}/{\mathbb{A}}^{\times}$ (see §2.4.1). Here is the basic identity, which is originally due to Andrianov in the classical setting ([1], [2]) and is adelically reformulated by Piatetski-Shapiro ([30]). To determine constants exactly under our normalization of Haar measures, we reproduce the proof briefly.
@@
×$\times\int_{{\mathbb{A}}^{\times}}\int_{{\bf K}_{\infty}^{\#}\,{\bf K}_{0}^{\#}(\mathfrak{o}_{E})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(k^{\#})\mu(a)|a|_{\mathbb{A}}^{s-1}\,B^{T_{\theta},\Lambda}(\varphi;\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#})b_{\mathbb{R}}^{\theta})\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}.$
@@
$L(s+1,\Lambda\mu_{E})^{-1}\times\langle E(\phi,s,\Lambda,\mu),R(b_{\mathbb{R}}^{\theta})\varphi\rangle$
$=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}})\mathsf{G}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}\sum_{\gamma\in{\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{Q}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(\gamma g^{\#})\,\varphi(\iota_{\theta}(g^{\#})b_{\mathbb{R}}^{\theta})\,{\rm{d}}g^{\#}$
$=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}}){\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(g^{\#})\,\varphi(\iota_{\theta}(g^{\#})b_{\mathbb{R}}^{\theta})\,{\rm{d}}g^{\#}$
@@ 4. Automorphic forms and Fourier coefficients
$\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(ng_{\infty}g_{\rm{\bf{f}}})\,\psi_{T}(n)^{-1}{\rm{d}}n$ $=\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n^{-1}\gamma_{\infty}g_{\infty}\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}})\,\psi_{T}(n)\,{\rm{d}}n$
$=\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n\gamma_{\infty}g_{\infty}\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}})\,\psi_{-T}(n)^{-1}\,{\rm{d}}n.$
@@
By the modular transformation law $f((AZ+B)(CZ+D)^{-1})=\varrho(CZ+D)^{-1}f(Z)$ for $\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\Gamma_{0}(N)$, since $-1_{4}\in\Gamma_{0}(N)$, we have that $f(Z)$ is identically 00, hence $S_{\varrho}({\bf K}_{0}(N))=(0)$, unless $l_{1}\equiv l_{2}\pmod{2}$. (For more details, see [37, §3.2] and [4, §4].)
@@
${\mathcal{Q}}:=\{\left[\begin{smallmatrix}a&b/2\\ b/2&c\end{smallmatrix}\right]\mid a,b,c\in{\mathbb{Z}}\}$
@@
$\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n)\,\psi_{-T{\mathsf{s}}\left(\delta\right)}(n)^{-1}\,{\rm{d}}n$ $=\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\mathsf{n}(X))\,\psi({\operatorname{tr}}(-(T{\mathsf{s}}\left(\delta\right))\,X))^{-1}\,{\rm{d}}X$
$=\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\mathsf{n}(-{\mathsf{s}}\left(\delta\right)X))\,\psi({\operatorname{tr}}(TX))^{-1}\,{\rm{d}}X$
@@
Define
(4.10) ${\mathcal{Q}}^{+}_{\rm prim}(D):=\{T=\left[\begin{smallmatrix}a&2^{-1}b\\ 2^{-1}b&c\end{smallmatrix}\right]\in{\mathcal{Q}}\mid a>0,\,-\det T=D/4\}.$
Since $D$ is a fundamental discriminant, $T\in{\mathcal{Q}}^{+}_{\rm prim}(D)$ is primitive, i.e., ${\gcd}(a,b,c)=1$; ${\mathcal{Q}}^{+}_{\rm prim}(D)$ is a subset of ${\mathsf{V}}({\mathbb{Q}})$, which is stable under the action of the unimodular group ${\bf SL}_{2}({\mathbb{Z}})$ induced by (2.1). The matrix $T_{\theta}$ defined by (2.11) belongs to ${\mathcal{Q}}_{\rm prim}^{+}(D)$. The ${\bf SL}_{2}({\mathbb{Z}})$-orbit of $T\in{\mathcal{Q}}^{+}_{\rm prim}(D)$ is denoted by $[T]$.
@@
For each $u\in{\mathbb{A}}_{E}^{\times}$, let $I_{\theta}(u)$ decompose in ${\bf GL}_{2}({\mathbb{A}})$ as
(4.11) $I_{\theta}(u)=\gamma_{u}h_{u}\kappa_{u},\quad\gamma_{u}\in{\bf GL}_{2}({\mathbb{Q}}),\,h_{u}\in{\bf GL}_{2}({\mathbb{R}})^{\circ},\,\kappa_{u}=(\kappa_{u,p})_{p<\infty}\in{\bf GL}_{2}(\widehat{\mathbb{Z}})$
@@
    $T_{\theta}(u)\in{\mathcal{Q}}_{\rm prim}^{+}(D)$.
@@
    The ${\bf SL}_{2}({\mathbb{Z}})$-equivalence class of $T_{\theta}(u)$ does not depend on the decomposition (4.11). If $u$, $u^{\prime}$ belongs to the same $E^{\times}E^{\times}_{\infty}\widehat{\mathfrak{o}}_{E}^{\times}$-coset, then $T_{\theta}(u)$ and $T_{\theta}(u^{\prime})$ are ${\bf SL}_{2}({\mathbb{Z}})$-equivalent.
@@
    The map $[u]\mapsto[T_{\theta}(u)]$ from ${\rm Cl}(E)={\mathbb{A}}_{E}^{\times}/E^{\times}E_{\infty}^{\times}\widehat{\mathfrak{o}}_{E}^{\times}$ to ${\mathcal{Q}}_{\rm prim}^{+}(D)/{\bf SL}_{2}({\mathbb{Z}})$ is a bijection.

### v2

@@ 2.4. Haar measures
We fix $\eta_{{\bf K}_{\infty}}$ so that ${\operatorname{vol}}({\bf K}_{\infty})=1$. Then we normalize $\eta_{{\bf Sp}_{2}({\mathbb{R}}))}$ in such a way that the quotient $\eta_{{\bf Sp}_{2}({\mathbb{R}})}/\eta_{{\bf K}_{\infty}}$ corresponds to the measure $(\det Z)^{-3}{\rm{d}}X{\rm{d}}Y$ on ${\mathfrak{h}}_{2}\cong{\bf Sp}_{2}({\mathbb{R}})/{\bf K}_{\infty}$. By $\mathsf{G}({\mathbb{R}})^{0}={\mathsf{Z}}({\mathbb{R}})^{0}\,{\bf Sp}_{2}({\mathbb{R}})\cong{\mathbb{R}}_{>0}\times{\bf Sp}_{2}({\mathbb{R}})$, $\eta_{\mathsf{G}({\mathbb{R}})}$ is defined by demanding that its restriction to $\mathsf{G}({\mathbb{R}})^{0}$ is $\eta_{{\mathbb{R}}^{\times}}\otimes\eta_{{\bf Sp}_{2}({\mathbb{R}})}$. For $p<\infty$, we fix $\eta_{\mathsf{G}({\mathbb{Q}}_{p})}$ by ${\operatorname{vol}}({\bf K}_{p})=1$. Then, $\eta_{\mathsf{G}({\mathbb{A}})}$ is defined to be the restricted product of $\eta_{\mathsf{G}({\mathbb{Q}}_{v})}\,(v\leqslant\infty)$.
2.5. Ray class characters
For each place $v$ of $E$, let ${\mathfrak{p}}_{v}$ be the corresponding maximal ideal of ${\mathfrak{o}}_{E}$. For any invertible ideal ${\mathfrak{m}}$ of ${\mathfrak{o}}_{E}$, define $U({\mathfrak{m}}):=\prod_{v<\infty}(1+{\mathfrak{p}}_{v}^{m_{v}}\mathfrak{o}_{E,v})\cap{\mathfrak{o}}_{E,v}^{\times}$ with $m_{v}:={\rm ord}_{v}{\mathfrak{m}}$, which is an open compact subgroup of $\widehat{\mathfrak{o}}_{E}^{\times}$. Let $\Lambda=\otimes_{v}\Lambda_{v}$ be an idele class character on ${\mathbb{A}}_{E}^{\times}$ of finite order such that $\Lambda|{\mathbb{A}}^{\times}={\bf 1}$; then $\Lambda|E_{\infty}^{\times}={\bf 1}$ and $\Lambda$ factors through a finite quotient
(2.22) ${\rm Cl}_{{\mathfrak{m}}}^{0}(E):={\mathbb{A}}_{E}^{\times}/{\mathbb{A}}^{\times}E^{\times}E_{\mathbb{R}}^{\times}U({\mathfrak{m}})$
for some invertible ideal ${\mathfrak{m}}$; the conductor of $\Lambda$ is the maximal one among such ideals ${\mathfrak{m}}$. As is well known, the group (2.22) is isomorphic to a quotient group of the ray class group of conductor ${\mathfrak{m}}$. We say that a place $v$ is inert in $E/{\mathbb{Q}}$, splits in $E/{\mathbb{Q}}$ or ramifies in $E/{\mathbb{Q}}$ if $E_{v}:=E\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ is an unramified field extension of ${\mathbb{Q}}_{p}$, is isomorphic to ${\mathbb{Q}}_{p}\oplus{\mathbb{Q}}_{p}$, or is a ramified field extension of ${\mathbb{Q}}_{p}$, respectively. We have the following:
[…]
Indeed, let $p$ be a prime below $v$; then $\Lambda_{v}(p)=1$ due to $\Lambda|{\mathbb{A}}^{\times}={\bf 1}$. Hence, $\Lambda_{v}$ is trivial on $p^{{\mathbb{Z}}}\mathfrak{o}_{E,v}^{\times}=E_{p}^{\times}$. Define the Galois conjugate $\Lambda^{\dagger}$ of $\Lambda$ by setting
@@ 2.5. Ray class characters
We have $\Lambda(\tau)\Lambda^{\dagger}(\tau)=\Lambda(\tau\bar{\tau})=1$ for $\tau\in{\mathbb{A}}_{E}^{\times}$ by $\Lambda|{\mathbb{A}}^{\times}={\bf 1}$. Hence
@@
For an idele class character $\xi$ of $E^{\times}$, let $\widehat{L}(s,\xi)$ be the completed Hecke $L$-function of $\xi$, and $L_{p}(s,\xi)$ its local $p$-factor for $p\leqslant\infty$. Then, $L_{p}(s,\xi^{\dagger})=L_{p}(s,\xi)$ for any $p<\infty$. Indeed, if $p$ is not ramified in $E/{\mathbb{Q}}$, the equality is trivial. Suppose that $p$ ramifies; if $\xi_{p}$ is a ramified character of $E_{p}^{\times}$, then both $L$-factors are $1$. If $\xi_{p}$ is unramified, then for any prime element $\varpi$ of $E_{p}$, we have $(\xi^{\dagger})_{p}(\varpi)=\xi_{p}(\bar{\varpi})=\xi_{p}(\varpi)$, which implies the equality between local $p$-factors.
For any character $\mu:{\mathbb{A}}^{\times}/{\mathbb{Q}}^{\times}{\mathbb{R}}_{>0}\rightarrow{\mathbb{C}}^{1}$, we define $\mu_{E}:=\mu\circ{\operatorname{N}}_{E/{\mathbb{Q}}}:{\mathbb{A}}^{\times}_{E}/E^{\times}\rightarrow{\mathbb{C}}^{1}$. Then, $\mu_{E}$ is Galois invariant, and $\mu_{E}|E_{\infty}^{\times}={\bf 1}$. Hence, $(\Lambda\mu_{E})^{\dagger}=\Lambda^{\dagger}\mu_{E}^{\dagger}=\Lambda^{-1}\mu_{E}$, and
@@ 3.1. Eisenstein series
The Fourier transform $\widehat{\phi}$ of $\phi\in\mathcal{S}(E_{p}^{2})$ is defined by
@@
with $c_{p}=1$ if $p<\infty$ and $c_{\infty}=\frac{\sqrt{|D|}}{2}$ if $p=\infty$, where $\varepsilon(s,\bullet,\psi_{E_{p}})$ denote Tate’s local epsilon factor defined by the character $\psi_{E_{p}}$ of $E_{p}$ and the associated self-dual measure on $E_{p}$ as usual. Let $\mathcal{S}({\mathbb{A}}_{E}^{2})$ be the space of all Schwartz-Buruhat functions on ${\mathbb{A}}_{E}^{2}$. For a decomposable element $\phi=\otimes_{p{\leqslant}\infty}\phi_{p}\in\mathcal{S}({\mathbb{A}}_{E}^{2})$, we define
@@
Note that $\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}$ is left ${\mathsf{B}}^{\#}({\mathbb{Q}})$- invariant and right ${\bf K}^{\#}$-finite. The Eisenstein series attached to $\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}$ is defined by the absolutely convergent series
@@
for ${\operatorname{Re}}(s)>1$. The Fourier transform $\widehat{\phi}$ of $\phi\in{\mathcal{S}}({\mathbb{A}}_{E}^{2})$ is defined by a formula similar to (3.3) with respect to the measure $\eta_{{\mathbb{A}}_{E}}\otimes\eta_{{\mathbb{A}}_{E}}$. The following property of the Eisenstein series is standard.
@@ 3.2. Rankin-Selberg integral and the basic identity
By Proposition 3.1(iii) and by the Fourier expansion (4.5), it is straightforward to show the absolute convergence of the integral for $s\in{\mathbb{C}}$ where the Eisenstein series is regular. For $T\in{\mathsf{V}}({\mathbb{Q}})$, define a character of $\mathsf{N}({\mathbb{A}})$ by
@@
The $(T_{\theta},\Lambda)$-Bessel period of a cusp form $\varphi$ on ${\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})$ is defined by the integral111 Note that $\mathsf{m}(I_{\theta}(a),{\operatorname{N}}_{E/{\mathbb{Q}}}(a))=a1_{4}$ for $a\in{\mathbb{A}}^{\times}$. Then, due to $\varphi$ being ${\mathsf{Z}}({\mathbb{A}})$-invariant, the integral $B^{T_{\theta},\Lambda}(\varphi;g)$ is 00 if $\Lambda|{\mathbb{A}}^{\times}\not=1$. This trivial vanishing is not the case by $\Lambda|{\mathbb{A}}^{\times}={\bf 1}$.
@@
for $g\in\mathsf{G}({\mathbb{A}})$, where ${\rm{d}}^{\times}\tau$ denote the quotient measure on ${\mathbb{A}}_{E}^{\times}/{\mathbb{A}}^{\times}$ (see §2.4.1). Here is the basic identity due to Piatetski-Shapiro ([31]). To determine constants exactly under our normalization of Haar measures, we reproduce the proof briefly.
@@
×$\times\int_{{\mathbb{A}}^{\times}}\int_{{\bf K}_{\infty}^{\#}\,{\bf K}_{0}^{\#}(\mathfrak{o}_{E})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(k^{\#})\mu(a)|a|_{\mathbb{A}}^{s-1}\,B^{T_{\theta},\Lambda}(\varphi;\mathsf{m}(a1_{2},a)\iota_{\theta}(k^{\#}))\,{\rm{d}}^{\times}a\,{\rm{d}}k^{\#}.$
@@
$L(s+1,\Lambda\mu_{E})^{-1}\times\langle E(\phi,s,\Lambda,\mu),\varphi\rangle$
$=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}})\mathsf{G}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}\sum_{\gamma\in{\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{Q}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(\gamma g^{\#})\,\varphi(\iota_{\theta}(g^{\#}))\,{\rm{d}}g^{\#}$
$=\int_{{\mathsf{Z}}^{\#}({\mathbb{A}}){\mathsf{B}}^{\#}({\mathbb{Q}})\backslash\mathsf{G}^{\#}({\mathbb{A}})}\mathsf{f}_{\phi}^{(s,\Lambda,\mu)}(g^{\#})\,\varphi(\iota_{\theta}(g^{\#}))\,{\rm{d}}g^{\#}$
@@ 4. Automorphic forms and Fourier coefficients
$\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(ng_{\infty}g_{\rm{\bf{f}}})\,\psi_{T}(n)^{-1}{\rm{d}}n$ $=\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n\gamma_{\infty}g_{\infty}\gamma_{\rm{\bf{f}}}g_{\rm{\bf{f}}})\,\psi_{-T}(n)^{-1}\,{\rm{d}}n.$
@@
By the modular transformation law $f((AZ+B)(CZ+D)^{-1})=\varrho(CZ+D)^{-1}f(Z)$ for $\left[\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right]\in\Gamma_{0}(N)$, since $-1_{4}\in\Gamma_{0}(N)$, we have that $f(Z)$ is identically 00, hence $S_{\varrho}({\bf K}_{0}(N))=(0)$, unless $l_{1}\equiv l_{2}\pmod{2}$. (For more details, see [38, §3.2] and [4, §4].)
@@
${\mathcal{Q}}:=\Bigl{\{}\left[\begin{smallmatrix}a&b/2\\ b/2&c\end{smallmatrix}\right]\,\Bigm{|}\,a,b,c\in{\mathbb{Z}}\Bigr{\}}$
@@
$\int_{\mathsf{N}({\mathbb{Q}})\backslash\mathsf{N}({\mathbb{A}})}\varphi(n)\,\psi_{-T{\mathsf{s}}\left(\delta\right)}(n)^{-1}\,{\rm{d}}n$ $=\int_{{\mathsf{V}}({\mathbb{Q}})\backslash{\mathsf{V}}({\mathbb{A}})}\varphi(\mathsf{n}(-{\mathsf{s}}\left(\delta\right)X))\,\psi({\operatorname{tr}}(TX))^{-1}\,{\rm{d}}X$
@@
For a positive integer ${\mathcal{M}}$, define
(4.10) ${\mathcal{Q}}^{+}_{{\mathcal{M}}}(D):=\{T=\left[\begin{smallmatrix}a&2^{-1}b\\ 2^{-1}b&c\end{smallmatrix}\right]\in{\mathcal{Q}}\mid a>0,\,{\gcd}(a,{\mathcal{M}})=1,\,-\det T=D/4\},$
${\bf\Gamma}_{0}({\mathcal{M}}{\mathbb{Z}}_{p}):={\bf SL}_{2}({\mathbb{Z}}_{p})\cap\left[\begin{smallmatrix}{\mathbb{Z}}_{p}&{\mathbb{Z}}_{p}\\ {{\mathcal{M}}}{\mathbb{Z}}_{p}&{\mathbb{Z}}_{p}\end{smallmatrix}\right]$, ${\bf\Gamma}_{0}({\mathcal{M}}\widehat{{\mathbb{Z}}}):=\prod_{p}{\bf\Gamma}_{0}({\mathcal{M}}{\mathbb{Z}}_{p})$, and $^{\Gamma}_{0}({\mathcal{M}}):={\bf SL}_{2}({\mathbb{Q}})\cap{\bf\Gamma}_{0}({\mathcal{M}}\widehat{\mathbb{Z}})$. Since $D$ is a fundamental, ${\mathcal{Q}}^{+}_{{\mathcal{M}}}(D)$ entirely consists of primitive forms, i.e., ${\rm gcd}(a,b,c)=1$; ${\mathcal{Q}}^{+}_{{\mathcal{M}}}(D)\subset{\mathsf{V}}({\mathbb{Q}})$ is ${\bf\Gamma}_{0}({\mathcal{M}})$-stable by the action (2.1); the ${\Gamma}_{0}({\mathcal{M}})$-orbit of $T\in{\mathcal{Q}}^{+}_{{\mathcal{M}}}(D)$ is denoted by $[T]$. Note that $T_{\theta}\in{\mathcal{Q}}_{{\mathcal{M}}}^{+}(D)$ (see (2.11)). Recall the notation in §2.5. When all the prime divisors of ${\mathcal{M}}$ are inert in $E/{\mathbb{Q}}$, we write $U({\mathcal{M}}):=U({\mathcal{M}}{\mathfrak{o}}_{E})$ and ${\rm Cl}_{{\mathcal{M}}}^{0}:={\rm Cl}_{{\mathcal{M}}{\mathfrak{o}}_{E}}^{0}$ briefly. Let $\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{{\rm prim}}$ be the set of all characters $\Lambda$ on ${\rm Cl}_{{\mathcal{M}}}^{0}(E)$ with conductor ${\mathcal{M}}{\mathfrak{o}}_{E}$
@@
Let ${\mathcal{M}}$ be a positive integer with prime factors being inert in $E/{\mathbb{Q}}$ and $\Lambda\in\widehat{{\rm Cl}_{{\mathcal{M}}}^{0}(E)}_{{\rm prim}}$. For each $u\in{\mathbb{A}}_{E}^{\times}$, let $I_{\theta}(u)$ decompose in ${\bf GL}_{2}({\mathbb{A}})$ as
(4.11) $I_{\theta}(u)=\gamma_{u}h_{u}\kappa_{u},\quad\gamma_{u}\in{\bf GL}_{2}({\mathbb{Q}}),\,h_{u}\in{\bf GL}_{2}({\mathbb{R}})^{\circ},\,\kappa_{u}=(\kappa_{u,p})_{p<\infty}\in{\bf\Gamma}_{0}({\mathcal{M}}\widehat{{\mathbb{Z}}})$
@@
    $T_{\theta}(u)\in{\mathcal{Q}}_{{\mathcal{M}}}^{+}(D)$.
@@
    The ${\Gamma}_{0}({\mathcal{M}})$-equivalence class of $T_{\theta}(u)$ does not depend on the decomposition (4.11). If $u$, $u^{\prime}$ belongs to the same ${\mathbb{A}}^{\times}E^{\times}E^{\times}_{\infty}U({\mathcal{M}})$-coset, then $T_{\theta}(u)$ and $T_{\theta}(u^{\prime})$ are ${\bf\Gamma}_{0}({\mathcal{M}})$-equivalent.
@@
    The map $[u]\mapsto[T_{\theta}(u)]$ from ${\rm Cl}_{{\mathcal{M}}}^{0}(E)$ to ${\mathcal{Q}}_{{\mathcal{M}}}^{+}(D)/{\Gamma}_{0}({\mathcal{M}})$ is a bijection.